OpenStudy (anonymous):

combining like terms... -2x + 11+ 6x =-8x is this correct

OpenStudy (turingtest):

-2x + 11+ 6x =-2x+6x+11 =(-2+6)x+11 =4x+11

OpenStudy (anonymous):

ty i wasnt even close